OpenStudy (anonymous):

solve the equation 9=square root of 3 ^(4x+6)Using complete sentences, explain the strategy used in solving this equation.

OpenStudy (amistre64):

those online courses are soo nitpicky arent they :)

OpenStudy (anonymous):

yes:(

OpenStudy (amistre64):

hard to parse the equation ..... for me that is

OpenStudy (anonymous):

9=¨3^2

OpenStudy (anonymous):

If you go back and look at the other 2 questions u posted like this, the way to do it will become clear...

OpenStudy (anonymous):

9=square root of 3 ^(4x+6) on squaring both sides => square of 9 = square of [ square root of 3 ^(4x+6) ] => 81 = 3^(4x+6) => 3^4 = 3^(4x+6) => 4 = 4x+6 //since a^x = a^y .=> x=y => 4x+6 = 4 => 4x = 4-6 => 4x = -2 => x = -2/4 // dividing with 4 on both sides => x = -1/2 => x = -0.5

OpenStudy (anonymous):

thank u!!!!!!!!!

OpenStudy (anonymous):

OMG...

OpenStudy (anonymous):

okay i get it now!

OpenStudy (anonymous):

3^2 = 3^(4x+6) SO 2 = 4x +6 -> x = -1

OpenStudy (anonymous):

i like the way u explained it and put it out thanks!

OpenStudy (anonymous):

u r welcome